OpenStudy (anonymous):

solve the system by substitution please help {2x+y=-11 {3x-4y=11 A. (3,5) B. (-5,-3) C. (-3,-5) D. (5,3)

OpenStudy (anonymous):

2x+y=-11------->y= -2x-11-----(1 {3x-4y=11-------(2 from 1 and 2 3x-4(-2x-11)=11 i.e. 3x+8x+44=11 i.e. 11x=11-44 i.e. 11x=-33 i.e. x=-33/11 i.e. x=-3 from eq 2 y= -2x-11 = -2(-3) -11=6-11= -5 Thus {x,y}={-3,-5} which is the required ans

OpenStudy (nurali):

2x+y=-11---(1) 3x-4y=11---(2) in equation (1) 2x+y=-11 2x=-11-y x=-11-y/2 put in equation (2) 3(-11-y/2)-4y=11 -33-3y/2-4y=11 -33-3y-8y=22 -33-11y=22 -11y=22+33 -11y=55 y=-55/11 y=-5 put in equation (1) 2x-5=-11 2x=-11+5 2x=-6 x=-6/2 x=-3 (x,y)=(-3,-5)